OpenStudy (anonymous):

greg is hiking through yosemite national park, but he didnt follow the marked trail. greg thinks he may be lost and must radio for help. using the data for his path find the compass cordinates of his location realtive to his starting position of (0,0) 15m East 5m South 3m North 1m West 13m West 21m east 20m west

OpenStudy (unklerhaukus):

 total displacement east = east -west \[=(15+21)\text{m}-(1+13+20)\text{m}\]\[=36\text{m}-34\text{m} \]\[=2\text mEast\]

OpenStudy (unklerhaukus):

now find the north south displacement north-south=\[(3-5)\text mNorth\] \[=(5-3)\text m South\]

OpenStudy (unklerhaukus):